# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ VI-E5 Sparse AEs
These SAEs impose sparsity constraints on hidden layers, ensuring that only a few neurones are activated at a time. This forces the AE to learn efficient and compact representations of the input data, even if the dimensionality of the latent space is not explicitly reduced. they are typically enforced by adding a regularisation term such as penalty to the loss function. This encourages the network to activate only a subset of neurones.
VI-E6 Adversarial AEs
@@
These AAEs utilise generative adversarial networks (GANs) to regularise the learnt latent space, making it match a specific prior distribution [42]. The goal is to achieve better generative capabilities, producing more realistic and diverse data compared to traditional AEs. Along with the reconstruction loss function, adversarial loss is used to ensure that latent space follows a target prior distribution. Compared to VAEs which use a probabilistic approach with KL divergence, AEs use adversarial training through a GAN framework.
@@
Future research should investigate in more depth the following AE architectures for the area of communication systems [42]:
  • Bayesian AEs
  • Diffusion AEs
  • Contractive AEs
  • Denoising AEs
  • Semi‑supervised AEs
  • Masked AEs
VII Geometric and Probabilistic Constellation Shaping
@@
Geometric and probabilistic constellation shaping are two techniques used to optimise the performance of communication systems [43, 44, 45, 46, 47]. Although they share some similarities, there are also significant differences between them. Table IV presents the main differences between them.
@@
GCS aims to optimise the constellation shape for better performance by optimising the spatial distribution of modulation symbols in a two-dimensional plane. That is, it involves designing a specific geometric shape for the modulation format to achieve better performance in terms of maximising the signal-to-noise ratio (SNR) or minimising the BER. This is typically done by defining a set of rules or constraints that define the geometry of the constellation and then optimising the modulation format accordingly. GCS can be achieved through various methods, such as mathematical optimisation techniques, graph theory-based approaches, or machine learning algorithms.
@@
On the other hand, probabilistic constellation shaping (PCS) involves controlling the probability distribution of constellation points. That is, it considers the geometric distribution and introduces a probabilistic aspect to it. Control both symbol distance and their occurrence probabilities in order to optimise system performance further. The aim is typically two-fold: to enhance spectral efficiency while simultaneously improving the SNR for better BER. This is typically done by optimising the modulation format to achieve better performance in terms of data rate, reach, and noise tolerance. Probabilistic shaping can be achieved through various methods such as optimisation algorithms, machine learning techniques, or signal processing approaches.
@@
Another key difference is the level of complexity involved in each approach. GCS can be simpler to implement, but may have limitations in terms of flexibility and adaptability to changing channel conditions. On the other hand, PCS can be more complex to implement and optimise due to its dual consideration - optimising both spatial and temporal aspects (occurrence probabilities), and it requires advanced mathematical techniques and signal processing algorithms. In the context of end-to-end learning based on AE in the communication system, joint geometric and probabilistic shaping can optimise the constellation for enhanced performance, such as maximisation of mutual information [48, 49, 50]. Fig 5 illustrates the two inputs to the mapper on the transmitter sides: the constellation points $c$ and their corresponding probabilities $p_{c}$. Optimising $c$, GCS is achieved, while optimising $p_{c}$ enables probabilistic constellation shaping.
TABLE IV: Geometric Constellation Shaping vs. Probabilistic Constellation Shaping
@@
optimise the spatial distribution of
@@
Performance Metrics maximise SNR, minimise BER Enhance spectral efficiency, improve SNR for better BER
@@ Methods
Mathematical optimisation, graph theory,
@@
optimisation algorithms, machine learning,
@@
Constellation Geometry optimised for specific geometries
Not optimised for specific geometries,
@@
VIII (non-) and Differentiable Channel and Components
A differentiable channel refers to a channel model that can be represented mathematically and whose behaviour can be differentiated with respect to its input. Specifically, it means that the channel output (received signal) can be expressed as a differentiable function of its input (transmitted signal). This property is crucial for training end-to-end neural network-based AE learning, where a single neural network represents the entire communication system (transmitter, channel, and receiver). The AE aims to learn an efficient representation of transmitted information, and having a differentiable channel allows gradients to flow through the entire system during training, as shown in Fig. 6.
VIII-A Determining Differentiability
To determining whether a channel is differentiable depends on the specific characteristics of the channel model. a differentiable function (channel) typically has a tangent line at every point along its graph. If the tangent exists, the function is continuous, and its derivative exists, making it differentiable [51]. Here are some considerations:
  • Mathematical formulation: If the channel can be described by a mathematical function, such as linear channel, AWGN, or Rayleigh fading, it is likely to be differentiable. For example, the AWGN channel can be expressed as $y=x+n$, where $y$ is the received signal, $x$ is the transmitted signal, and $n$ is the noise. This channel is differentiable because its behaviour can be expressed as a continuous function.
  • Discrete Effects: Non-differentiability often arises from discrete effects, such as quantisation and modulation. If the channel introduces discrete operations, such as a threshold, it may not be differentiable. In such cases, alternative training methods such as reinforcement learning are needed.
  • Noise Models: Noise models matter. AWGN is differentiable due to its Gaussian noise, but other noise models, such as impulsive noise, may not be.
  • Physical limitations: Real-world channels have physical constraints, such as power limits or bandwidth limits. These constraints affect differentiability. For instance, clipping can make a channel non-differentiable.
VIII-B Solution Approaches for Non-differentiable
  • Alternating training algorithm: The idea behind this is to optimise the transmitter and receiver independently. The iterative optimisation process involves alternating between optimising the receiver with a fixed transmitter and subsequently optimising the transmitter with a fixed receiver. This approach aims to enhance the overall system performance through this iterative refinement. This can be achieved through an alternating optimisation approach, where the receiver is trained using supervised learning while the transmitter is optimised using reinforcement learning [53, 54].
  • Differentiable DSP algorithms: An end-to-end differentiable AE channel require all components, including DSP algorithms, to be differentiable to enable gradient-based optimisation. Traditional DSP algorithms often incorporate non-differentiable operations, hindering their direct integration into the differentiable channel. Therefore the authors of [55] proposed to modify blind phase search (BPS) algorithm and make it differentiable to include it in the end-to-end constellation shaping.
  • Learning stochastic channel model: This approach involves learning a surrogate channel model, either through supervised learning [56] or adversarial methods [57] The key idea is to learn a differentiable generative model of the channel in the form of a GAN, which can then be used to train the AE, and to approximate end-to-end channel responses. The learned model is subsequently employed to train the transceiver. This enables direct optimisation of the modulation and encoding components within an end-to-end AE model. The accuracy of this learned model is crucial for the overall system performance. In [58], a channel-sensitive AE (CSAE) incorporates a pre-trained conditional GAN (CGAN) to model optical fibre communication systems with varying impairments.
  • Two-phase training strategy [52]: With this approach, first train the AE using a channel model designed closely to the real channel such as a stochastic channel model. Upon deploying the trained model for real-wold transmission, performance is contingent on the model’s accuracy. To mitigate this, based on measurement data a find-tuning phase focusing exclusively on the receiver part of the AE, which lead to a sub-optimal training.
  • Reinforcement learning-based approaches: the idea is to train an neural network (NN)-based transmitter using policy gradients while employing a non-differentiable receiver that treats detection as a clustering problem [59].
  • Stochastic perturbation techniques: A gradient free optimisation techniques called Simultaneous perturbation stochastic optimisation has been proposed in [60]. it is a novel, single-stage approach to train deep learning-based communication systems directly in real-world channel environments. By circumventing the need for explicit channel modelling, this way leverages stochastic approximation to estimate gradients, enabling end-to-end optimisation.
  • Gradient-free training method: It is a derivative-free optimisation method to eliminate the need for a differentiable channel model in AE training. This enables simultaneous optimisation of the encoder and decoder for arbitrary, even non-numerical, channel environments. The authors of [61] proposed the use of the cubature kalman filter (CKF) which offers gradient-free training while maintaining accuracy, thereby facilitating online adaptation.
    Non-differentiableChannel SolutionsAlternatingtraining algorithmLearning stochasticchannel modelTwo-phase training strategyReinforcementlearning-based approachesDiffusion ModelsGenetic algorithmsGradient-freetraining methodDifferentiableDSP algorithmsStochastic PerturbationTechnique
    Figure 7: Solutions for Non-differentiable Channel
  • Diffusion models: The idea is to the use of diffusion models (DMs) to accurately learn channel distributions. The work in [62] focusses on generative models and proposes an efficient training algorithm for end-to-end coded-modulation systems. DMs have demonstrated high sample quality and mode coverage in image generation, where the resulting differentiable channel model supports training neural encoders by enabling gradient-based optimisation.
  • Genetic algorithms (GAs): While traditional end-to-end learning relies heavily on backpropagation, which requires the system to be differentiable, genetic algorithms offer an alternative optimisation method that does not rely on gradient information [63, 64, 65, 66].
    • optimisation without gradients: GAs optimise a population of solutions through evolutionary processes such as selection, crossover, and mutation. This optimisation process does not require the computation of gradients and can be applied to non-differentiable functions.
    • Fitness evaluation: In the context of an AE, the GAs can evaluate the fitness of each individual in the population based on how well the AE reconstructs the input data after it has passed through the channel. The fitness function can be designed to measure the reconstruction error or any other relevant performance metric.
    • Non-differentiable channels: If the channel in the communication system is non-differentiable, due to involving quantisation, discrete noise, or other non-continuous processes, traditional gradient-based methods fail to optimise the AE effectively. GAs, on the other hand, can still optimise the system by treating it as a black-box and relying on fitness evaluations rather than gradient information.
    • Population diversity: GAs maintain a population of potential solutions, which can help explore the solution space more thoroughly compared to gradient descent methods that might get stuck in local minima. This can be particularly beneficial in complex, non-differentiable landscapes.
IX AE Complexity
@@
Computational complexity is essential for both the encoder and decoder components of AE-based systems. A practical approach involves counting floating-point operations (FLOPs) [67], which is commonly used to assess the computational performance of algorithms and hardware in machine learning. It measures the computational performance of algorithms and hardware, which is more practical for machine learning algorithms. Additionally, counting FLOPs for algorithm complexity has different advantages such as efficiency analysis of the algorithm, hardware benchmarking, and optimising the algorithm. In the context of AEs, FLOPs must be calculated for both the encoder and the decoder. For each layer, the number of FLOPs for matrix multiplication, bias addition and nonlinear activation functions should be considered as follows:
IX-A Matrix Multiplication
@@
The AE encoder receives a one-hot vector $1_{l}\in\{0,1\}^{L}$ of length $L$. This input is then multiplied by a weight matrix $W\in R^{N\times L}$, where $N$ is the number of neurones in the layer. Following [67], the computational complexity of matrix multiplication between matrices $X$ and $Y$ with sizes ($a\times b$) and ($b\times c$) respectively, is given by ($2abc-ac$). Therefore, in the context of an AE layer with $N$ neurones, the number of FLOPs required is $N(2L-1)$.
IX-B Bias Addition
@@
Each neurone is associated with bias addition represented by a vector $B\in R^{L}$, where $L$ is the number of neurones in a layer. Adding the bias term requires one FLOP per neurone. Therefore, the total number of FLOPs for both matrix multiplication and bias addition can be calculated as $N(2L-1)+L$.
IX-C Nonlinear Activation Function
@@
Table  V provides a breakdown of the FLOPs counted for various activation functions in a layer with $N$ neurones. Appendix  A details the specific calculations for each activation function.

### v2

@@ V-E5 Sparse AEs
These SAEs impose sparsity constraints on hidden layers, ensuring that only a few neurons are activated at a time. This forces the AE to learn efficient and compact representations of the input data, even if the dimensionality of the latent space is not explicitly reduced. they are typically enforced by adding a regularization term such as penalty to the loss function. This encourages the network to activate only a subset of neurons.
V-E6 Adversarial AEs
@@
These AAEs utilize generative adversarial networks (GANs) to regularize the learned latent space, making it match a specific prior distribution [51]. The goal is to achieve better generative capabilities, producing more realistic and diverse data compared to traditional AEs. Along with the reconstruction loss function, adversarial loss is used to ensure that latent space follows a target prior distribution. Compared to VAEs which use a probabilistic approach with KL divergence, AEs use adversarial training through a GAN framework.
@@
Future research should investigate in more depth the following AE architectures for the area of communication systems [51]: Bayesian AEs, Diffusion AEs, Contractive AEs, Denoising AEs, Semi‑supervised AEs, and Masked AEs
V-F Geometric and Probabilistic Constellation Shaping
@@
Geometric and probabilistic constellation shaping are two techniques used to optimize the performance of communication systems [52, 53, 54, 55, 56]. Although they share some similarities, there are also significant differences between them. Table III presents the main differences between them.
@@
GCS aims to optimize the constellation shape for better performance by optimizing the spatial distribution of modulation symbols in a two-dimensional plane. That is, it involves designing a specific geometric shape for the modulation format to achieve better performance in terms of maximizing the signal-to-noise ratio (SNR) or minimizing the BER. This is typically done by defining a set of rules or constraints that define the geometry of the constellation and then optimizing the modulation format accordingly. GCS can be achieved through various methods, such as mathematical optimization techniques, graph theory-based approaches, or machine learning algorithms.
@@
On the other hand, probabilistic constellation shaping (PCS) involves controlling the probability distribution of constellation points. That is, it considers the geometric distribution and introduces a probabilistic aspect to it. Control both symbol distance and their occurrence probabilities in order to optimize system performance further. The aim is typically two-fold: to enhance spectral efficiency while simultaneously improving the SNR for better BER. This is typically done by optimizing the modulation format to achieve better performance in terms of data rate, reach, and noise tolerance. Probabilistic shaping can be achieved through various methods such as optimization algorithms, machine learning techniques, or signal processing approaches.
@@
Another key difference is the level of complexity involved in each approach. GCS can be simpler to implement, but may have limitations in terms of flexibility and adaptability to changing channel conditions. On the other hand, PCS can be more complex to implement and optimize due to its dual consideration, optimizing both spatial and temporal aspects (occurrence probabilities), and it requires advanced mathematical techniques and signal processing algorithms. In the context of end-to-end learning based on AE in the communication system, joint geometric and probabilistic shaping can optimize the constellation for enhanced performance, such as maximization of mutual information [57, 58, 59]. Fig 5 illustrates the two inputs to the mapper on the transmitter sides: the constellation points $c$ and their corresponding probabilities $p_{c}$. Optimizing $c$, GCS is achieved, while optimizing $p_{c}$ enables probabilistic constellation shaping.
TABLE III: Geometric Constellation Shaping vs. Probabilistic Constellation Shaping
@@
optimize the spatial distribution of
@@
Performance Metrics maximize SNR, minimize BER Enhance spectral efficiency, improve SNR for better BER
@@ Methods
Mathematical optimization, graph theory,
@@
optimization algorithms, machine learning,
@@
Constellation Geometry optimized for specific geometries
Not optimized for specific geometries,
@@
V-G AE Complexity
This section briefly outlines the calculation methodology of the AE complexity model. An examination of the literature on the complexity of AE is presented in Section VI-F1. Moreover, Section IX-3 offers a discussion of future developments concerning AE complexity and computational constraints.
@@ V-G AE Complexity
Computational complexity is essential for both encoder and decoder components of AE-based systems. A practical approach involves counting floating-point operations (FLOPs) [60], which is commonly used to assess the computational performance of algorithms and hardware in machine learning. It measures the computational performance of algorithms and hardware, which is more practical for machine learning algorithms. Additionally, counting FLOPs for algorithm complexity has different advantages such as efficiency analysis of the algorithm, hardware benchmarking, and optimizing the algorithm. In the context of AEs, FLOPs must be calculated for both the encoder and the decoder. For each layer, the number of FLOPs for matrix multiplication, bias addition, and nonlinear activation functions should be considered as follows:
V-G1 Matrix Multiplication
@@
The AE encoder receives a one-hot vector $1_{l}\in\{0,1\}^{L}$ of length $L$. This input is then multiplied by a weight matrix $W\in R^{N\times L}$, where $N$ is the number of neurons in the layer. Following [60], the computational complexity of matrix multiplication between matrices $X$ and $Y$ with sizes ($a\times b$) and ($b\times c$) respectively, is given by (2abcac2abc-ac). Therefore, in the context of an AE layer with $N$ neurons, the number of FLOPs required is N(2L1)N(2L-1).
V-G2 Bias Addition
@@
Each neuron is associated with bias addition represented by a vector $B\in R^{L}$, where $L$ is the number of neurons in a layer. Adding the bias term requires one FLOP per neuron. Therefore, the total number of FLOPs for both matrix multiplication and bias addition can be calculated as N(2L1)+LN(2L-1)+L.
V-G3 Nonlinear Activation Function
@@
Table IV provides a breakdown of the FLOPs counted for various activation functions in a layer with $N$ neurons.